# Exploring Graph-Transformer Out-of-Distribution Generalization Abilities

### v1

@@ Abstract
In this work, we address the challenge of out-of-distribution (OOD) generalization for graph neural networks, with a special focus on the impact of backbone architecture. We systematically evaluate GT and hybrid backbones in OOD settings and compare them to MPNNs. To do so, we adapt several leading domain generalization (DG) algorithms to work with GTs and assess their performance on a benchmark designed to test a variety of distribution shifts. Our results reveal that GT and hybrid GT-MPNN backbones consistently demonstrate stronger generalization ability compared to MPNNs, even without specialized DG algorithms.
Additionally, we propose a novel post-training analysis approach that compares the clustering structure of the entire ID and OOD test datasets, specifically examining domain alignment and class separation. Demonstrating its model-agnostic design, this approach not only provided meaningful insights into GT and MPNN backbones. It also shows promise for broader applicability to DG problems beyond graph learning, offering a deeper perspective on generalization abilities that goes beyond standard accuracy metrics. Together, our findings highlight the promise of graph-transformers for robust, real-world graph learning and set a new direction for future research in OOD generalization.
@@ 1 Introduction
Robust evaluation under OOD conditions is thus crucial for deploying models in high-stakes applications. However, this evaluation is challenging, as traditional i.i.d benchmarks are inadequate and require dedicated datasets and metrics to reflect realistic distribution shifts. The Graph Out-of-Distribution (GOOD) benchmark [8] addresses this by providing diverse data splits with various distribution shifts. Still, it reports results solely on a single backbone, virtual-GIN (vGIN) [3, 9], with DG algorithms trained on top to enhance OOD robustness. While a few recent works have begun to explore the impact of different backbones on OOD generalization [10, 11], they remain limited, focusing only on MPNN variants or node-level tasks, often excluding strong-performing baselines like GIN. These observations highlight the need to study how backbone architectures affect OOD performance independently of DG algorithms, and whether modern designs like transformer-based backbones might improve robustness to distribution shifts.
@@
Besides the exploration and influence of the backbone of OOD performance, another key open question in the study of generalization abilities is how to evaluate and interpret OOD performance effectively. Most existing work relies primarily on accuracy-based metrics or intuitively inspecting features visualization [20]. In this work, we question whether accuracy alone is sufficient to evaluate OOD generalization and instead propose quantifying decoupled aspects that capture deeper model behavior. This issue is especially relevant in high-stakes applications, where selecting models that perform reliably under distribution shifts is critical. To this end, we employ post-hoc interpretability (analyzing a trained model without altering its architecture) by loading a model checkpoint and propose the examination of its internal representations on the full in-distribution (ID) and OOD test datasets.
@@
When considering post-hoc interpretability, a key step is defining which model outputs or representations should be evaluated. Softmax outputs are commonly used for classification, yet they are known to be unreliable and overconfident for inputs that lie outside the training distribution [21]. The penultimate layer is usually used for visualization, including in OOD contexts [22, 23], yielding higher-dimensional features, which tend to yield more robust uncertainty estimates. Beyond selecting the right representations, proper evaluation also demands dedicated metrics that capture decoupled aspects of generalization. Haotian Ye et al. [24] introduce the concepts of variation and informativeness as critical components for evaluating a model’s ability to generalize to OOD data. Variation refers to the change in learned features across different distributions, while informativeness measures how well a feature distinguishes between classes. Misalignment in these embeddings often signals a failure to generalize effectively to OOD inputs [25].
@@
To move beyond the limitations of analyzing individual samples, we suggest a more global perspective on generalization evaluation. While some OOD detection methods effectively identify anomalies by comparing OOD sample features to ID training features alignment [26], they operate at the individual-level. In contrast, since OOD generalization aims to assess a model’s overall robustness, we argue that examining the global structure of both ID and OOD datasets provides deeper insight. To this end, we propose a cluster-wise analysis that compares the overall clustering structure of model representations across domains. First, we propose employing the squared maximum mean discrepancy (MMD) [27] distance to measure alignment between ID and OOD features clusters, in a post-hoc manner. Li et al. [28] also use MMD for OOD generalization, but as a training loss to align source domain features and learn domain-invariant representations. Second, we propose to use the Silhouette score [29], which evaluates how well each sample fits within its own class compared to other classes, capturing cluster cohesion and separation. We suggest applying it to concatenated ID and OOD features using class labels, which reveals how well the class structure is preserved across domains and whether OOD samples overlap or disrupt ID clusters. This approach enables us to reveal patterns that are invisible to sample-wise comparisons, offering a more interpretable and comprehensive quantification of model robustness under distribution shift.
@@ 1.1 Contributions
  • We conduct the first systematic evaluation of GT backbones, including both pure GT and hybrid GT-MPNN (GPS) backbones, for OOD generalization in supervised graph classification. This includes adapting several DG algorithms to support GT backbones, and rigorously comparing them to an MPNN backbone (vGIN) across four diverse distribution shifts using the GOOD benchmark.
@@
CMNIST-colorMotif-sizeMotif-basisSST2-length0020202020404040406060606080808080100100100100OOD AccuracyvGINMHAGPS
Figure 1: OOD accuracy in % of three backbones across four benchmarks. vGIN is equipped with DG algorithms (CORAL, DANN), while MHA and GPS are trained using standard ERM. See Section A for experimental details.
Figure 2: Overview of the GOOD benchmark graph classification tasks. Key properties and visual examples highlight distribution shifts across CMNIST, Motif, and SST2 tasks. Figure adapted from S. Gui et al. [8].
@@ 2 Problem definition
This work focuses on OOD generalization in graph classification tasks. Our problem can be defined as DG that relies on incorporating domain-shift information. Formally, we consider a dataset with a set of $|\mathcal{E}|$ domains =$\mathcal{E}=\{e_{1},...,e_{|E|}\}$. Samples $(G^{e}_{i},Y^{e}_{i})\in e$ from the same domain are considered to have been drawn independently from an identical distribution $\mathbb{P}^{e}(G,Y)$. Let $\mathcal{E}_{all}$ be the domain set we want to generalize to, and the training and test domains $\mathcal{E}_{train},\mathcal{E}_{test}\subset\mathcal{E}_{all}$. The goal of OOD generalization on graphs is to train a classifier $f^{*}$ with data from training domains $\mathcal{E}_{train}$ that generalizes well to all domains $\mathcal{E}_{all}$:
[…]
@@ 3 Methodology
This section describes our key components and analysis framework for tackling OOD generalization on graphs. We first review the GPS layer, a hybrid architecture combining local message passing and global graph transformer attention for richer graph representations. Next, we discuss the strengths and limitations of GT and MPNN backbones for OOD tasks. We then review several DG approaches aimed at improving robustness to unseen data distributions and employed in our experiments. Finally, to quantitatively assess OOD generalization in the latent feature space, we introduce two metrics that evaluate its quality: MMD, which measures distributional shifts between ID and OOD embeddings, and the Silhouette score, which evaluates the clarity and separation of clusters formed by these representations. Together, they provide deeper and decoupled insights about the model’s generalization abilities, given a specific dataset (or task).
@@ 3.2 Graph Transformer as a backbone for OOD tasks
There are several reasons in favor of GT. 1) Global context modeling is crucial for large graphs and tasks where the predicted label is casual to the entire or a large portion of the graph, as message-passing is computationally feasible only for two or three hops. 2) Multi-head attention (MHA) with flexible receptive field, which allows focusing on relevant graph parts, while ignoring irrelevant ones. Even though the MPNN variant, graph attention networks [34], may also be thought to possess those abilities, keep in mind that it only has a local receptive field. 3) MPNN’s over-smoothing and over-squashing can lead to information loss, compromise substructural integrity for larger diameter substructures, and insufficient edge information extraction. GT mitigate these issues by aggregating information over all nodes in a graph, avoiding local structure bias and ensuring accurate node-edge relationships. 4) As shown by [14], GT are more expressive than MPNNs bounded by 1-WL, because they can approximate any function using precomputed positional encodings based on the eigen decomposition of the graph Laplacian (universal approximation theorem [35, 36].)
GT, while offering advantages in certain scenarios, might not always be the best choice for OOD tasks compared to MPNNs for several reasons. 1) The over-globalizing problem [37] in GT arises when nearby nodes hold most useful information. In such cases, local pattern modeling is crucial, while attention may overemphasize distant nodes (though GPS, with its hybrid global-local modeling, may address this). 2) The attention mechanism requires significant computational resources, particularly when dealing with large graphs (though it’s a well-studied topic with some "sparse" solutions). 3) Virtual nodes [9] for message-passing might be a simple alternative, facilitating global information exchange. The above advantages and limitations of GT have been studied in the i.i.d context, but lack an in-depth evaluation for OOD.
@@ 3.3 Domain generalization algorithms
Data augmentation approaches generate more diverse training data, leading to better OOD generalization ability [38]. These approaches include structure-wise augmentations (which potentially cover some unobserved testing topologies), feature-wise augmentations (by manipulating node features), and a combination of both. Specifically, Manifold-Mixup [39] manipulates the training distribution in the semantic space by generating virtual graph representations.
@@
Learning strategy approaches concentrate on utilizing training schemes with specific optimization objectives and constraints to promote OOD generalization. Invariant learning aims to learn features or representations that are invariant across different domains during training [42]. Particularly, researchers have actively pursued feature distribution alignment using methods such as the MMD [28], and second-order correlation [43], among others [44, 45, 46]. Adversarial training to minimize domain discrepancy in the embedding space. Adversarial learning focuses on reducing the distribution discrepancy in the embedding space across different domains [47].
Note that in Section 4, we show that in certain cases the backbone architecture choice is even more critical than DG algorithms, as it determines feature extraction quality and significantly affects overall model performance.
@@ 3.4 Squared Maximum Mean Discrepancy (MMD)
To quantify alignment between domain distributions, we employ the squared MMD [27] distance in our generalization abilities analysis in Section 4.5. This process involves loading a trained model checkpoint, processing both ID and OOD test datasets to extract penultimate layer features, and then calculating the domain alignment using squared MMD.
We compute the MMD distance between two features sets $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{j}\}_{j=1}^{m}$ using a radial basis function (RBF) kernel based on the L1 distance:
@@ 3.5 Silhouette score
To quantify the quality of clustering in a feature representation, we use the Silhouette score [29] in our generalization abilities analysis in Section 4.5. This process involves loading a trained model checkpoint, processing both ID and OOD test datasets to extract penultimate layer features, and then calculating the class-separation using the Silhouette score.
@@ 4.1 Experimental setup
Experiments are conducted on four graph-classification datasets from the GOOD benchmark: CMNIST (color shift), Motif (size shift), Motif (basis shift), and SST2 (length shift). For each dataset, covariate shift splits are applied based on domain features to create diverse distribution shifts between the training set, OOD validation set, and OOD test set. The training set is then shuffled and further divided into a final training set, ID validation set, and ID test set. Figure 2 summarizes their key properties, such as input types, prediction tasks, and domain selection, and visually shows how training and test graph structures differ, illustrating the distribution shifts.
@@
For the backbone architectures, we use vGIN, MHA, and GPS. Our baseline is vGIN, an MPNN variant that extends GIN with a virtual node to enhance global information flow, consistent with the backbone choice in the GOOD benchmark. For the pure GT backbone, we use MHA, which applies a global attention mechanism between graph nodes. We selected random-walk (RW) embeddings [48] as positional encodings, based on their demonstrated effectiveness on ID tasks [49]. For the hybrid MPNN-GT backbone, we use GPS, as described in Section 3.1. Note that GPS uses standard GIN (not vGIN) for local message passing and MHA-RW for the global receptive field. All models are trained using the hyperparameters listed in Table 3, with the number of trainable parameters matched across backbones (per-task) as closely as possible to ensure a fair comparison. Global mean pooling is applied for node-level aggregation. Additional details are provided in Appendix A.
@@ 4.2 Evaluation metrics
While OOD accuracy (measured on the OOD test set) is the most prevalent metric for quantifying OOD performance, it alone does not capture all aspects of generalization such as accuracy degradation between the ID to OOD settings. To address this, we also report the ID-OOD accuracy gap, defined as $gap=ID_{test}\text{-}OOD_{test}$, which is a common metric that is used in NLP and computer-vision [50, 18]. Referred to simply as the accuracy gap throughout the paper, this metric reflects the model’s sensitivity to distribution shifts relative to its ID performance and complements the OOD accuracy metric.
To assess the model’s generalization capabilities beyond standard accuracy metrics, we evaluated the squared MMD distance (defined in Section 3.4) between the penultimate-layer representations of ID and OOD test datasets. This provides a quantitative measure of feature space alignment across domains. Furthermore, we computed the Silhouette score (see Section 3.5) on the concatenated penultimate-layer features from both ID and OOD datasets to quantify how well each data point fits within its assigned cluster (intra-cluster distance) compared to other clusters (inter-cluster distance).
We selected the best checkpoint, in our opinion, on the ID validation set to compute all four evaluation metrics, as this better reflects real-world scenarios where OOD data is not available during training or validation. Final evaluation is then conducted and reported on both ID and OOD test sets.

### v2

@@ Abstract
In this work, we address the challenge of out-of-distribution (OOD) generalization for graph neural networks, with a special focus on the impact of backbone architecture. We systematically evaluate GT and hybrid backbones in OOD settings and compare them to MPNNs. To do so, we adapt several leading domain generalization (DG) algorithms to work with GTs and assess their performance on a benchmark designed to test a variety of distribution shifts. Our results reveal that GT and hybrid GT-MPNN backbones demonstrate stronger generalization ability compared to MPNNs, even without specialized DG algorithms (on four out of six benchmarks).
Additionally, we propose a novel post-training analysis approach that compares the clustering structure of the entire ID and OOD test datasets, specifically examining domain alignment and class separation. Highlighting its model-agnostic design, the method yielded valuable insights into both GT and MPNN backbones and appears well suited for broader DG applications beyond graph learning, offering a deeper perspective on generalization abilities that goes beyond standard accuracy metrics. Together, our findings highlight the promise of graph-transformers for robust, real-world graph learning and set a new direction for future research in OOD generalization.
@@ 1 Introduction
Robust evaluation under OOD conditions is thus crucial for deploying models in high-stakes applications. However, this evaluation is challenging, as traditional i.i.d benchmarks are inadequate and require dedicated datasets and metrics to reflect realistic distribution shifts. The Graph Out-of-Distribution (GOOD) benchmark [8] addresses this by providing diverse data splits with various distribution shifts. Still, it reports results solely on a single backbone, virtual-GIN (vGIN) [3, 9], with DG algorithms trained on top to enhance OOD robustness. While a few recent works have begun to explore the impact of different backbones on OOD generalization [10, 11], they remain limited, focusing only on MPNN variants or node-level tasks, often excluding strong-performing baselines like GIN. Although our prior work was the first to investigate graph transformers in OOD settings, a recent concurrent and independent study, GOODFormer [12], proposed a DG algorithm for transformers that jointly optimizes invariant subgraph disentanglement and evolving subgraph encoding. Their focus was primarily on the DG algorithm rather than systematically evaluating the backbone itself, which resulted in the transformer-based backbones not achieving their full performance potential on most benchmarks. These observations emphasize the need to systematically study how backbone architectures influence OOD performance independently of DG algorithms, and whether modern designs, such as transformer-based backbones, can enhance robustness to distribution shifts.
@@
Besides the exploration and influence of the backbone of OOD performance, another key open question in the study of generalization abilities is how to evaluate and interpret OOD performance effectively. Most existing work relies primarily on accuracy-based metrics or intuitively inspecting features visualization [21]. In this work, we question whether accuracy alone is sufficient to evaluate OOD generalization and instead propose quantifying decoupled aspects that capture deeper model behavior. To this end, we employ post-hoc interpretability, analyzing a trained model without altering its architecture, by loading a model checkpoint and examining its internal representations across the full in-distribution (ID) and OOD test datasets.
@@
When considering post-hoc interpretability, an important step is defining which model representations should be evaluated. For classification tasks, softmax outputs are often used, but they can be unreliable and overconfident for inputs that lie outside the training distribution [22]. Higher-dimensional features from the penultimate layer are commonly used instead [23, 24], as they tend to provide more robust uncertainty estimates and richer information for analysis. Beyond selecting the right representations, proper evaluation also demands dedicated metrics that capture decoupled aspects of generalization. Haotian et al. [25] introduce the concepts of variation and informativeness as critical components for evaluating a model’s ability to generalize to OOD data. Variation refers to the change in learned features across different distributions, while informativeness measures how well a feature distinguishes between classes. Misalignment in these learned features often signals a failure to generalize effectively to OOD inputs [26].
@@
To move beyond the limitations of analyzing individual samples, we suggest a more global perspective on generalization evaluation. While some OOD detection methods effectively identify anomalies by comparing OOD sample features to ID training features alignment [27], they operate at the individual-level. In contrast, since OOD generalization aims to assess a model’s overall robustness, we argue that examining the global structure of both ID and OOD datasets provides deeper insight. To this end, we propose a cluster-wise analysis that compares the overall clustering structure of model representations across domains. First, we propose employing the squared maximum mean discrepancy (MMD) [28] distance to measure alignment between ID and OOD features clusters, in a post-hoc manner. Li et al. have used MMD for OOD generalization in a different way, as a training loss to align source-domain features and learn domain-invariant representations. Second, we propose to use the Silhouette score [29], which evaluates how well each sample fits within its own class compared to other classes, capturing cluster cohesion and separation. Applied to concatenated ID and OOD features with class labels, it reveals how well the class structure is preserved across domains and whether OOD samples overlap or disrupt ID clusters. This approach enables to reveal patterns that are invisible to sample-wise comparisons, offering a more interpretable and comprehensive quantification of model robustness under distribution shift.
@@ 1.1 Contributions
  • We conduct the first systematic evaluation of GT backbones, including both pure GT and hybrid GT-MPNN (GPS) backbones, for OOD generalization in supervised graph classification. This includes adapting several DG algorithms to support GT backbones, and rigorously comparing them to an MPNN backbone (vGIN) across six diverse distribution shifts using the GOOD benchmark.
@@
CMNIST-colorMotif-sizeMotif-basisSST2-lengthHIV-sizeHIV-scaffold02020404060608080100100OOD AccuracyvGINMHAGPS
Figure 1: OOD accuracy in % of three backbones across six benchmarks. vGIN is equipped with DG algorithms (which lead to the best OOD accuracy), while MHA and GPS are trained using standard ERM only. See Section A for experimental details.
Figure 2: Overview of the GOOD benchmark graph classification tasks. Key properties and visual examples highlight distribution shifts across CMNIST, Motif, SST2, and HIV tasks. Figure adapted from S. Gui et al. [8].
@@ 2 Problem definition
This work focuses on OOD generalization in graph classification tasks. The problem can be defined as DG that relies on incorporating domain-shift information. Formally, we consider a dataset with a set of $|\mathcal{E}|$ domains $\mathcal{E}=\{e_{1},...,e_{|E|}\}$. Samples $(G^{e}_{i},Y^{e}_{i})\in e$ from the same domain are considered to have been drawn independently from an identical distribution $\mathbb{P}^{e}(G,Y)$. Let $\mathcal{E}_{all}$ be the domain set we want to generalize to, and the training and test domains $\mathcal{E}_{train},\mathcal{E}_{test}\subset\mathcal{E}_{all}$. The goal of OOD generalization on graphs is to train a classifier $f^{*}$ with data from training domains $\mathcal{E}_{train}$ that generalizes well to all domains $\mathcal{E}_{all}$:
$f^{*}=\underset{f}{\arg\min}\underset{G,Y\in\mathbb{P}^{\mathcal{E}_{all}}}{\mathbb{E}}[\mathcal{L}(f(G),Y)].$ (1)
@@ 3 Methodology
This section describes our key components and analysis framework for tackling OOD generalization on graphs. We first review the GPS layer, a hybrid architecture combining local message passing and global graph transformer attention for richer graph representations. Next, we discuss the strengths and limitations of GT and MPNN backbones for OOD tasks. We then review several DG approaches aimed at improving robustness to unseen data distributions and employed in our experiments. Finally, to quantitatively assess OOD generalization in the latent feature space, we introduce two metrics that evaluate its quality. MMD, which measures distributional shifts between ID and OOD embeddings, and the Silhouette score, which evaluates the clarity and separation of clusters formed by these representations. Together, they provide deeper and decoupled insights about the model’s generalization abilities, given a specific dataset (or task).
@@ 3.2 Graph Transformer as a backbone for OOD tasks
There are several reasons in favor of GT. First, global context modeling is crucial for large graphs and tasks where the predicted label is casual to the entire or a large portion of the graph, as message-passing is computationally feasible only for two or three hops. Second, multi-head attention (MHA) with flexible receptive field, which allows focusing on relevant graph parts, while ignoring irrelevant ones. Even though the MPNN variant, graph attention networks [34], may also be thought to possess those abilities, keep in mind that it only has a local receptive field. Third, MPNN’s over-smoothing and over-squashing can lead to information loss, compromise substructural integrity for larger diameter substructures, and insufficient edge information extraction. GT mitigate these issues by aggregating information over all nodes in a graph, avoiding local structure bias and ensuring accurate node-edge relationships. Fourth, as shown by [15], GT are more expressive than MPNNs bounded by 1-WL, because they can approximate any function using precomputed positional encodings based on the eigen decomposition of the graph Laplacian (universal approximation theorem [35, 36].)
GT, while offering advantages in certain scenarios, might not always be the best choice for OOD tasks compared to MPNNs for several reasons. First, the over-globalizing problem [37] in GT arises when nearby nodes hold most useful information. In such cases, local pattern modeling is crucial, while attention may overemphasize distant nodes (though GPS, with its hybrid global-local modeling, may address this). Second, the attention mechanism requires significant computational resources, particularly when dealing with large graphs, though various “sparse” approaches have been proposed to mitigate this cost [38]). Third, virtual nodes [9] for message-passing might be a simple alternative, facilitating global information exchange. The above advantages and limitations of GT have been studied in the i.i.d context, but lack an in-depth evaluation for OOD.
@@ 3.3 Domain generalization algorithms
Data augmentation approaches generate more diverse training data, leading to better OOD generalization ability [39]. These approaches include structure-wise augmentations, which potentially cover some unobserved testing topologies, feature-wise augmentations (by manipulating node features), and a combination of both. Specifically, Manifold-Mixup [40] manipulates the training distribution in the semantic space by generating virtual graph representations.
@@
Learning strategy approaches focus on training schemes with specific objectives and constraints to promote OOD generalization. Invariant learning aims to extract features or representations that remain consistent across domains during training [43]. This often involves explicitly aligning feature distributions using techniques such as MMD [44], second-order correlation [45], or methods like IRM and others [46, 47, 48]. In contrast, adversarial learning enforces domain invariance indirectly by training the feature extractor to fool a domain discriminator [49].
In Section 4, we show that in certain cases the backbone architecture choice is even more critical than DG algorithms, as it determines feature extraction quality and significantly affects overall model performance.
@@ 3.4 Squared Maximum Mean Discrepancy (MMD)
To quantify alignment between domain distributions, the squared MMD [28] distance is employed for the analysis of generalization abilities in Section 4.5. This process involves loading a trained model checkpoint, processing both ID and OOD test datasets to extract penultimate layer features, and then calculating the domain alignment using squared MMD.
The MMD distance between two features sets $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{j}\}_{j=1}^{m}$ is computed using a radial basis function (RBF) kernel based on the L1 distance:
@@ 3.5 Silhouette score
The quality of clustering in a feature representation is quantified using the Silhouette score [29] for the analysis of generalization abilities in Section 4.5. This process involves loading a trained model checkpoint, processing both ID and OOD test datasets to extract penultimate layer features, and then calculating the class-separation using the Silhouette score.
@@ 4.1 Experimental setup
Experiments are conducted on six graph-classification datasets from the GOOD benchmark: CMNIST (color shift), Motif (size shift), Motif (basis shift), SST2 (length shift), HIV (size shift), and HIV (scaffold shift). For each dataset, covariate shift splits are applied based on domain features to create diverse distribution shifts between the training set, OOD validation set, and OOD test set. The training set is then shuffled and further divided into a final training set, ID validation set, and ID test set. Figure 2 summarizes their key properties, such as input types, prediction tasks, and domain selection, and visually shows how training and test graph structures differ, illustrating the distribution shifts.
@@
For the backbone architectures, we use vGIN, MHA, and GPS. Our baseline is vGIN, an MPNN variant that extends GIN with a virtual node to enhance global information flow, consistent with the backbone choice in the GOOD benchmark (except for Motif tasks, where we used GIN instead of vGIN). For the pure GT backbone, MHA, which applies a global attention mechanism between graph nodes, is used. We selected random-walk (RW) embeddings [50] as positional encodings, based on their demonstrated effectiveness on ID tasks [51]. For the hybrid MPNN-GT backbone, we use GPS, as described in Section 3.1. Note that GPS uses standard GIN, rather than vGIN, for local message passing and MHA-RW for the global receptive field. All models are trained using the hyperparameters listed in Table 3, with the number of trainable parameters matched across backbones (per-task) as closely as possible to ensure a fair comparison. Global mean pooling is applied for node-level aggregation. Additional details are provided in Appendix A.
@@ 4.2 Evaluation metrics
While OOD accuracy, measured on the OOD test set, is the most prevalent metric for quantifying OOD performance, it alone does not capture all aspects of generalization such as accuracy degradation between the in-distribution (ID) to OOD settings. To address this, we also report the ID-OOD accuracy gap, defined as $gap=ID_{test}\text{-}OOD_{test}$, which is a common metric that is used in NLP and computer-vision [52, 19]. Referred to simply as the accuracy gap throughout the paper, this metric reflects the model’s sensitivity to distribution shifts relative to its ID performance and complements the OOD accuracy metric.
To assess the model’s generalization capabilities beyond standard accuracy metrics, we evaluated the squared MMD distance between the penultimate-layer representations of ID and OOD test datasets. This provides a quantitative measure of feature space alignment across domains. Furthermore, we computed the Silhouette score on the concatenated penultimate-layer features from both ID and OOD datasets to quantify how well each data point fits within its assigned cluster (intra-cluster distance) compared to other clusters (inter-cluster distance).
The best checkpoint is selected based on the ID validation set, reflecting realistic scenarios where OOD data is unavailable during training or validation. This contrasts with prior works that select checkpoints using an OOD validation set, inflating performance beyond realistic deployment conditions. Final evaluation is then conducted on both ID and OOD test sets.